# DiffGAN: A Test Generation Approach for Differential Testing of Deep Neural Networks for Image Analysis

### v3

@@ 3.3.1 Individuals and Initial Population
In genetic search, the initial population contains individuals consisting of a set of elements called genes. These genes are connected and form an individual also referred to as a solution. In our context, the initial population contains 100 randomly generated latent vectors, each of size 100, as recommended in the literature [36]. Each dimension of a latent vector in the initial population is independently generated from a uniform distribution over the range [1,1][-1,1], thereby ensuring a diverse and unbiased initialization in the search space.
@@ 3.3.2 Fitness Functions
This fitness function estimates the divergence between the output probability vectors of the models under test, when fed with the same input image generated by the trained GAN. More specifically, the fitness function takes as input a GAN latent vector $Z$ responsible for generating the candidate test input image. We feed the models under test with such image and compute the Euclidean distance between the resulting output probability vectors $Y_{m1}$ and $Y_{m2}$. The goal of maximizing this distance is to increase the probability of selecting latent vectors that lead to the generation of input images revealing significant behavioral disagreements between the models under test.
The Euclidean distance serves as a straightforward measure of the magnitude of discrepancy between model predictions, capturing how far apart the models’ outputs are in the probability space. The choice of Euclidean distance over other metrics, such as KL divergence [37], is motivated by its efficiency in terms of computation cost.
@@
However, the use of Euclidean distance or any other distance metrics alone has a limitation: they do not account for the differences in predicted labels across models. This means that while a high Euclidean distance indicates a discrepancy in the output probabilities, it does not necessarily reflect a disagreement in the models’ final predicted labels. This is particularly critical when the primary objective lies in searching for triggering inputs that lead to different label predictions by the models under test.
@@
We add the following example to illustrate the above limitation. Let us consider that for a generated image (image 1), the Euclidean distance calculated between the output probability vectors P1 = [0.05,0.05,0.05,0.85] from the first model under test (Model 1) and Q1 = [0.15,0.15,0.15,0.55] from the second model under test (Model 2) is 0.3. Further, both models predict the same class for image 1. Then, let us consider another image (image 2), where the calculated Euclidean distance between the output probability vectors P2 = [0.1,0.2,0.3,0.4] from Model 1 and Q2 = [0.15,0.25,0.35,0.05] from Model 2 is also 0.3, despite that these models predicting different classes for image 2. We obtain the same Euclidean distances for both images though clearly image 2 is more interesting given our objectives.
@@
Using a standard Euclidean distance in the context of differential testing is therefore not optimal and indicates the need to refine the fitness function, to ensure it not only considers the magnitude of input differences but also changes in the predicted labels, indicating different models predictions.
@@
To address this limitation, we modified the fitness function to reflect differences in model predictions. More specifically, given a GAN latent vector, if the models output different labels for a GAN-generated image, the fitness function increments the Euclidean distance by one to always assign a higher score for triggering input images. Otherwise, it returns the Euclidean distance between the two output probability vectors as illustrated in Equation 1:
@@
Given a latent vector, this fitness function computes the contribution of the corresponding GAN-generated image to the diversity of the collected triggering inputs during the search process. To do so, the diversity-based fitness function computes the minimum cosine distance score between the pixel values of this generated image and those of the set of collected triggering inputs during the search.
To reduce the computational complexity of comparing each generated image against all collected triggering inputs at each iteration of the search process, we approximate such distance by applying a clustering algorithm to group similar inputs. Then, we only return the minimum cosine distance between the generated image and the representative triggering inputs for all clusters. Preliminary experiments confirmed that, despite the cost of clustering, significant time could be saved in computing similarity scores of new images, thus leading to generating more triggering inputs. More specifically, at each iteration of the search process, we apply the HDBSCAN [38] clustering algorithm on the set of the collected triggering input images to group similar inputs in one cluster. We select HDBSCAN for its ability to identify clusters of varying densities without requiring the prior specification of the number of clusters. This adaptability makes HDBSCAN particularly suited in our context.
Prior to clustering, we implement UMAP for dimensionality reduction, which assists in forming more distinct clusters [39, 27] at a lower cost. Subsequently, within each cluster, we identify the representative input—specifically, the input that has the maximum probability of association with its corresponding cluster222 HDBSCAN Documentation—retrieved October 5, 2024, from https://hdbscan.readthedocs.io/en/latest/soft_clustering_explanation.html. The fitness function then calculates the diversity of the GAN-generated image by measuring the cosine distances between this image and the representatives of all clusters. Finally, the diversity-based fitness function returns the minimum of these cosine distances to quantify the diversity contribution of the image, as illustrated in Equation 2:
@@ 3.4.1 Crossover
The crossover operator takes as input two parents (i.e., two latent vectors) and generates new offspring by combining the selected parents. Similar to existing works that employ NSGA-II on GAN latent vectors [28, 40], we use the simulated binary crossover (SBX) in our approach as it is suitable for real-coded individuals (i.e., continuous optimization problem) [40, 41, 28]. Unlike traditional binary crossover operators, which are designed for discrete or binary representations, SBX operates directly on the real values of the latent vectors. Let $V_{1}$ and $V_{2}$ be the two selected parents for crossover. We provide next an example of the resulting offspring after applying crossover:
@@ 3.4.2 Mutation
Our genetic search also employs a mutation operator to further explore the search space.
Following existing works that use genetic algorithms in GAN latent space [28, 40], we also adopt polynomial mutation [42] to address the continuous nature of the latent space.
Polynomial mutation applies polynomial-distributed noise to the genes forming the selected individuals in our search. By applying such noise to the genes, it allows for the exploration of nearby solutions while still maintaining proximity to the original solution. This is also essential for efficiently navigating the complex and high-dimensional search space of GAN latent vectors. Polynomial mutation helps generate diverse solutions by introducing variability in the population of the search. This diversity is crucial for escaping local optima and finding globally optimal or near-optimal solutions.
@@ 21 $gen=gen+1$;
The main objective of our search algorithm is to maximize the generation of diverse triggering inputs to effectively identify the discrepancies between the DNN models under test. A high-level overview of our NSGA-II search process is outlined in Algorithm 1. Assuming $P$ represents the initial population, consisting of a set of $|P|$ random latent vectors, the algorithm initiates an iterative procedure, executing the following steps in each generation until the maximum execution time ($T$) is reached (lines 3-14). The search process includes the following steps. First, we create a new empty population $P_{new}$ (line 4). Second, based on predefined crossover and mutation rates, we create offspring using crossover ($Cross$) and mutation ($Mutate$) operators, incorporating newly created individuals to $P_{new}$ (lines 5-7). Then, for each individual, we generate the corresponding image with the trained GAN and store the triggering images in an archive (lines 8-12). Next, we calculate the fitness scores of the new population by computing the diversity- and divergence-based fitness scores of each individual in the population. We employ tournament [43] selection ($Select$) to choose individuals for survival (line 13). Lastly, we update the population by incorporating the fittest individuals, proceeding to the next generation (line 14). The search terminates when the maximum execution time ($T$) is reached and returns the archive of generated triggering inputs (line 15).
We set the NSGA-II parameters in our search algorithm as follows. The crossover rate CrossRateCrossRate is equal to 90% and the mutation rate MutationRateMutationRate is equal to 10%, given recommended values in the literature [44, 45, 46]. Finally, we used Google Colab and the Pymoo library [47] to implement our genetic search.
@@ 4.1 Research Questions
We aim to study in this research question the effectiveness of DiffGAN in producing more triggering inputs compared to a baseline, namely DRfuzz, within the same testing budget. Our choice of DRfuzz is carefully justified below, along with a description of the technique. The testing budget in our context is defined by a predetermined and fixed execution time. Generating large numbers of triggering inputs that reveal behavioral disagreement between DNN models is essential to accurately learn when each of the models can be expected to fare better. Relying on too few triggering images would not enable DiffGAN to identify the subtle differences between models, especially those with similar, high accuracy. Therefore, our research question aims to evaluate and contrast the efficiency of DiffGAN against DRfuzz in terms of their ability to generate more triggering inputs within time constraints.
@@
Generating large numbers of triggering inputs is not sufficient. To the largest extent possible, they also need to be valid. We, therefore, aim to compare the validity of the triggering inputs produced by DiffGAN and those generated by DRfuzz. The concept of input validity is crucial for DNN testing in general and differential testing in particular since it ensures that the generated inputs are both relevant and meaningful [3, 10]. More specifically, they should be valid and belong to the targeted input domain. A valid input for a DNN model is defined as one that human experts within the domain can recognize and categorize accurately, using labels from the domain [3]. For example, in the context of handwritten digit recognition, an image is considered valid if it can be recognized as a digit by a human expert, despite potential imperfections. We, therefore, aim to compare in this research question the proportions of valid inputs that DiffGAN and DRfuzz generate. This comparison is crucial because generating a large number of triggering inputs is only beneficial if these inputs are to a large extent valid, as defined above.
@@
We aim in this research question to determine whether the triggering inputs generated by DiffGAN can more effectively guide the training of a model selection mechanism, allowing it to learn the conditions under which one DNN model outperforms another. In many real-world scenarios, it is common to encounter multiple DNNs that perform similarly in terms of overall accuracy, but exhibit significant variations in their performance under specific input conditions [51, 52, 10, 11]. This presents a challenge for developing a model selection mechanism that can dynamically choose the most suitable model based on the input characteristics without having to run all models in real-time. An effective solution to this problem is to use an ML-based selection mechanism, which can predict which of the available models will perform better for a given input [51, 52]. However, training such an ML-based model selection mechanism requires diverse and informative training data that captures and reveals a wide range of behavioural differences between the models. This is where the generation of triggering inputs becomes critical for the selection mechanism to more effectively learn under which condition one model is preferable over the other. Indeed, the effectiveness of the selection mechanism heavily depends on the quality of the triggering inputs used for training. Therefore, we aim to explore in this RQ whether DiffGAN can better guide the training process compared to DRfuzz. It is important to clarify that we do not intend to propose a novel ensemble method or voting mechanism in this work. Instead, this research question focuses on investigating an important application scenario, that is studying whether DiffGAN helps learn which model fares better under various input conditions and thus train a model selection mechanism.
@@ 4.3 Datasets and Models
We consider different combinations of datasets and models in our empirical evaluation. More specifically, we consider MNIST [54] and Cifar-10 [55], two well-known image recognition datasets. As shown in Figure 2, the MNIST dataset consists of 70,000 black-and-white images depicting handwritten digits, split into a training set of 60,000 images and a test set of 10,000 images. The Cifar-10 dataset comprises 60,000 color images distributed across ten categories (Figure 3), including 50,000 images designated for training and 10,000 images for testing. These datasets were employed in conjunction with 16 different DNN models, which included both custom-developed and widely used architectures such as ResNet, LeNet-5, AlexNet, and inception models [56, 57, 58]. Comprehensive details about the combinations of datasets and models we used, along with their accuracy, are outlined Tables I and II.
As shown in the tables, our experimental study spanned four distinct and common sources of variation per dataset across DNN models, including the model structures, training data, optimizers and hyperparameters. Our experiments thus cover 16 different DNN models, each selected to evaluate the effectiveness of our test generation approach. For each model pair, the differences in DNN models are highlighted in bold in Tables I and II. Typical differences across models include:
@@ 4.4 Choice of the Fitness Functions
Due to the computational expense of our experiments, we conducted this study on two model pairs with the Cifar-10 dataset. Seeing significant differences across configurations on such pairs would however be enough to justify the combined use of divergence and diversity. We studied both the total number of triggering inputs generated by each configuration and their diversity. To quantify diversity, we leveraged two metrics: Shannon Diversity [59, 60] and Geometric Diversity [1]. A detailed explanation of these metrics, along with our methodology for computing diversity scores, is provided in Section 5.3. The results of this ablation study are summarized in Table III.
@@ 5 Empirical Results
We describe in this section our experimental evaluation and report on the results. We first investigate whether we generate more triggering inputs than the baseline with the same testing budget. We then study the validity and diversity of the generated triggering inputs by both approaches. Finally, we present an important application scenario of the generated triggering inputs, namely the development of a machine learning-based mechanism to predict the best model’s output based on input characteristics, and report the corresponding results.
@@
To answer this research question, we executed DiffGAN and DRfuzz considering the same execution time $T\in[1h,2h,3h]$, across four pairs of DNN models and two distinct datasets, resulting in eight unique combinations (Section 4.3). Because of the randomness in DiffGAN and DRfuzz, we re-executed each of them three times and reported the corresponding number of triggering inputs in Tables IV and V. This decision to limit the execution of the approaches to three iterations was initially due to the considerable execution time associated with the experiments. More specifically, our experiments take 96h (8 model pairs x 6h x 2 approaches) if we execute each approach once. As shown in the tables, DiffGAN consistently outperforms DRfuzz in terms of generating more triggering inputs across all models and datasets. More specifically, our results show that DiffGAN generates on average four times more triggering inputs than DRfuzz when considering the same testing budget. We also calculated the Coefficient of Variation (CV) [61], a measure of relative variability that compares the standard deviation to the mean, independent of the data’s scale. We computed the CV score for each model pair and testing budget across the three runs. A lower CV indicates higher consistency across different runs. For example, in our analysis of the MNIST dataset, DRfuzz showed a CV between 0.33% and 21.23%,while DiffGAN exhibited a CV between 1.33% and 12.20%, suggesting less variability in the number of generated triggering inputs across the different runs. These observations are also consistent for the Cifar-10 dataset. These CV values demonstrate the relative stability and reliability of DiffGAN in generating triggering images, with less variability compared to DRfuzz.
@@
Table V: RQ1 results: Number of generated triggering images for models trained on Cifar-10 dataset using DiffGAN and DRfuzz. The #TI (Initial) column indicates the number of triggering inputs identified in the original test dataset for each set of models.
@@
proach):
To systematically compare DiffGAN and DRfuzz in terms of generating valid triggering inputs within a given time budget, we considered the inputs generated in RQ1 by both methods. More specifically, we considered all four model pairs using the MNIST dataset and two model pairs for the CIFAR-10 dataset. We ran both approaches for two hours on these models. We did not consider all model pairs in this RQ as the validity assessments require significant manual effort, particularly for labeling generated inputs to ensure they are contextually valid. The lower resolution and greater complexity of CIFAR-10 images further complicate this process, as accurately labeling and validating inputs in CIFAR-10 is considerably more challenging compared to MNIST. Given these challenges, we opted to limit our validity study to the MNIST dataset to ensure its feasibility.
@@
Manual validity checking. From the non-redundant set, we randomly sampled images for validation, according to the sample size determined the previous step. To ensure unbiased evaluation, we shuffled images generated by both DiffGAN and DRfuzz and presented them anonymously to three independent expert evaluators. These experts were tasked with assessing the validity of each image based on predefined criteria. Specifically, they labeled images as either valid or invalid. For example for MNIST dataset, invalid images were defined as those that displayed two numbers, no number, or other issues such as (1) Images that were a 50/50 blend of two digits and could not be clearly identified as a specific digit, (2) Incomplete or blended digits that could not be classified into one category by consensus, (3) Images containing random shapes or non-digit drawings, or (4) Images that were completely white or dark, with no discernible content. Examples of valid and invalid inputs generated by both approaches are depicted in Figure 4. To determine the validity of each image, we considered the agreement among the evaluators. If at least two out of three evaluators labeled an image as valid, we considered it valid. To quantify the level of agreement between the evaluators, we employ Cohen’s kappa [63]—a statistical measure that accounts for agreement occurring by chance. In our study, the value of Cohen’s kappa is 70%, indicating a high level of agreement among the evaluators and underscoring the reliability of the manual labeling process.
@@
The metric computes the sum of the proportions of each category multiplied by the logarithm of those proportions. The negative sign ensures that the index is positive and higher diversity translates into higher values of $H$. To calculate the Shannon diversity for a set of images, each image is first flattened into a one-dimensional array of pixel values. These flattened arrays can then be combined into a single array that represents the distribution of pixel values across the entire image set. The Shannon score, or entropy, is then calculated using this distribution, quantifying the diversity of pixel intensities within the dataset. A higher Shannon score indicates greater diversity within the set of images.
In some contexts, computing the exponential of the Shannon diversity score is used to provide a more intuitive interpretation of the index. Specifically, by calculating $e^{H}$, where $H$ is the Shannon diversity of a set of images, we can derive the effective number of image categories within the set. This value represents the number of equally likely categories that would produce the same level of entropy as observed in the dataset. For example, if $H$ increases from 11 to 12, the exponential value $e^{12}$ is substantially larger (approximately 162,754.79) when compared to $e^{11}$ = 59,874.14, indicating that the dataset is considerably more diverse. This approach allows to translate entropy into a more intuitive measure of diversity, making it easier to understand and compare the diversity of different datasets.
@@
In this research question, we employed both the Shannon and Geometric diversity metrics to analyze the dataset, aiming to evaluate image diversity from two complementary perspectives. The Shannon diversity metric offers a probabilistic view, emphasizing the distribution of pixel values, while the Geometric diversity metric assesses diversity based on spatial relationships and geometric patterns within the image features extracted by VGG16. By integrating these two metrics, we achieve a more comprehensive understanding of the datasets’ diversity.
@@
Models Method Only Model A is correct Only Model B is correct Both models are incorrect
@@
When properly trained with triggering inputs, the RF model should learn when each classifier is more accurate, effectively serving as a decision mechanism for selecting the optimal classifier for a given input. We choose RF because: (1) it scales effectively to handle a large number of features, and (2) its robustness against overfitting [71, 72, 73]. We do not intend to provide a novel voting mechanism here but, instead, provide an example of an important application scenario for the triggering inputs we generate. It also serves as a basis to compare the value of the inputs generated by DiffGAN to enable such model selection and compare it with that of DRfuzz.

### v4

@@ 3.3.1 Individuals and Initial Population
In genetic search, the initial population contains individuals consisting of a set of elements called genes. These genes are connected and form an individual, also referred to as a solution. In our context, the initial population contains 100 randomly generated latent vectors, each of size 100, as recommended in the literature [36]. Each dimension of a latent vector in the initial population is independently generated from a uniform distribution over the range [1,1][-1,1], thereby ensuring a diverse and unbiased initialization in the search space.
@@ 3.3.2 Fitness Functions
This fitness function estimates the divergence between the output probability vectors of the models under test when fed with the same input image generated by the trained GAN. More specifically, the fitness function takes as input a GAN latent vector $Z$ responsible for generating the candidate test input image. We feed the models under test with such an image and compute the Euclidean distance between the resulting output probability vectors $Y_{m1}$ and $Y_{m2}$. The goal of maximizing this distance is to increase the probability of selecting latent vectors that lead to the generation of input images revealing significant behavioral disagreements between the models under test. The Euclidean distance serves as a straightforward measure of the magnitude of discrepancy between model predictions, capturing how far apart the models’ outputs are in the probability space. The choice of Euclidean distance over other metrics, such as KL divergence [37], is motivated by its efficiency in terms of computation cost.
@@
However, the use of Euclidean distance or any other distance metric alone has a limitation: they do not account for the differences in predicted labels across models. This means that while a high Euclidean distance indicates a discrepancy in the output probabilities, it does not necessarily reflect a disagreement in the models’ final predicted labels. This is particularly critical when the primary objective lies in searching for triggering inputs that lead to different label predictions by the models under test.
@@
We add the following example to illustrate the above limitation. Let us consider that for a generated image (image 1), the Euclidean distance calculated between the output probability vectors P1 = [0.05,0.05,0.05,0.85] from the first model under test (Model 1) and Q1 = [0.15,0.15,0.15,0.55] from the second model under test (Model 2) is 0.3. Further, both models predict the same class for image 1. Then, let us consider another image (image 2), where the calculated Euclidean distance between the output probability vectors P2 = [0.1,0.2,0.3,0.4] from Model 1 and Q2 = [0.15,0.25,0.35,0.05] from Model 2 is also 0.3, despite the fact that these models predict different classes for image 2. We obtain the same Euclidean distances for both images, though clearly image 2 is more interesting given our objectives.
@@
Using a standard Euclidean distance in the context of differential testing is therefore not optimal and indicates the need to refine the fitness function, to ensure it not only considers the magnitude of input differences but also changes in the predicted labels, indicating different models’ predictions.
@@
To address this limitation, we modified the fitness function to reflect differences in model predictions. More specifically, given a GAN latent vector, if the model’s output different labels for a GAN-generated image, the fitness function increments the Euclidean distance by one to always assign a higher score for triggering input images. Otherwise, it returns the Euclidean distance between the two output probability vectors as illustrated in Equation 1:
@@
Given a latent vector, this fitness function computes the contribution of the corresponding GAN-generated image to the diversity of the collected triggering inputs during the search process. To achieve this, the diversity-based fitness function calculates the minimum cosine distance score between the pixel values of the generated image and those of the set of collected triggering inputs during the search.
To reduce the computational complexity of comparing each generated image against all collected triggering inputs at each iteration of the search process, we approximate such distance by applying a clustering algorithm to group similar inputs. Then, we only return the minimum cosine distance between the generated image and the representative triggering inputs for all clusters. Preliminary experiments confirmed that, despite the cost of clustering, significant time could be saved in computing similarity scores of new images, thereby leading to the generation of more triggering inputs. More specifically, at each iteration of the search process, we apply the HDBSCAN [38] clustering algorithm on the set of the collected triggering input images to group similar inputs in one cluster. We select HDBSCAN for its ability to identify clusters of varying densities without requiring the prior specification of the number of clusters. This adaptability makes HDBSCAN particularly suited in our context.
Prior to clustering, we implement UMAP for dimensionality reduction, which assists in forming more distinct clusters [39, 27] at a lower cost. Subsequently, within each cluster, we identify the representative input—specifically, the input that has the maximum probability of association with its corresponding cluster333 HDBSCAN Documentation—retrieved October 5, 2024, from https://hdbscan.readthedocs.io/en/latest/soft_clustering_explanation.html. The fitness function then calculates the diversity of the GAN-generated image by measuring the cosine distances between this image and the representatives of all clusters. Finally, the diversity-based fitness function returns the minimum of these cosine distances to quantify the diversity contribution of the image, as illustrated in Equation 2:
@@ 3.4.1 Crossover
The crossover operator takes as input two parents (i.e., two latent vectors) and generates new offspring by combining the selected parents. Similar to existing works that employ NSGA-II on GAN latent vectors [28, 40], we use the simulated binary crossover (SBX) in our approach as it is suitable for real-coded individuals (i.e., continuous optimization problem) [40, 41, 28]. Unlike traditional binary crossover operators, which are designed for discrete or binary representations, SBX operates directly on the real values of the latent vectors. Let $V_{1}$ and $V_{2}$ be the two selected parents for crossover. We provide the following example of the resulting offspring after applying crossover:
@@ 3.4.2 Mutation
Our genetic search also employs a mutation operator to further explore the search space. Following existing works that use genetic algorithms in GAN latent space [28, 40], we also adopt polynomial mutation [42] to address the continuous nature of the latent space.
Polynomial mutation applies polynomial-distributed noise to the genes forming the selected individuals in our search. By applying such noise to the genes, it enables the exploration of nearby solutions while maintaining proximity to the original solution. This is also essential for efficiently navigating the complex and high-dimensional search space of GAN latent vectors. Polynomial mutation helps generate diverse solutions by introducing variability in the population of the search. This diversity is crucial for escaping local optima and finding globally optimal or near-optimal solutions.
@@ 21 $gen=gen+1$;
The main objective of our search algorithm is to maximize the generation of diverse triggering inputs to effectively identify the discrepancies between the DNN models under test. A high-level overview of our NSGA-II search process is outlined in Algorithm 1. Assuming $P$ represents the initial population, consisting of a set of $|P|$ random latent vectors, the algorithm initiates an iterative procedure, executing the following steps in each generation until the maximum execution time ($T$) is reached (lines 3-14). The search process includes the following steps. First, we create a new empty population $P_{new}$ (line 4). Second, based on predefined crossover and mutation rates, we create offspring using crossover ($Cross$) and mutation ($Mutate$) operators, incorporating newly created individuals to $P_{new}$ (lines 5-7). Then, for each individual, we generate the corresponding image with the trained GAN and store the triggering images in an archive (lines 8-12). Next, we calculate the fitness scores of the new population by computing the diversity- and divergence-based fitness scores of each individual in the population. We employ tournament selection (SelectSelect[43] to choose individuals for survival (line 13). Lastly, we update the population by incorporating the fittest individuals, proceeding to the next generation (line 14). The search terminates when the maximum execution time ($T$) is reached and returns the archive of generated triggering inputs (line 15).
We set the NSGA-II parameters in our search algorithm as follows. The crossover rate CrossRateCrossRate is set to 90% and the mutation rate MutationRateMutationRate is set to 10%, as recommended in the literature [44, 45, 46]. Finally, we used Google Colab and the Pymoo library [47] to implement our genetic search.
@@ 4.1 Research Questions
We aim to study in this research question the effectiveness of DiffGAN in producing more triggering inputs compared to a baseline, namely DRfuzz, within the same testing budget. Our choice of DRfuzz is carefully justified below, along with a description of the technique. The testing budget in our context is defined by a predetermined and fixed execution time. Generating large numbers of triggering inputs that reveal behavioral disagreement between DNN models is essential to accurately learn when each of the models can be expected to fare better. Relying on too few triggering images would not enable DiffGAN to identify the subtle differences between models, especially those with similar high accuracy. Therefore, our research question aims to evaluate and contrast the efficiency of DiffGAN against DRfuzz in terms of their ability to generate more triggering inputs within time constraints.
@@
Generating large numbers of triggering inputs is not sufficient. To the largest extent possible, they also need to be valid. We, therefore, aim to compare the validity of the triggering inputs produced by DiffGAN and those generated by DRfuzz. The concept of input validity is crucial for DNN testing in general and differential testing in particular since it ensures that the generated inputs are both relevant and meaningful [3, 10]. More specifically, they should be valid and belong to the targeted input domain. A valid input for a DNN model is defined as one that human experts within the domain can accurately recognize and categorize, using labels from the domain [3]. For example, in the context of handwritten digit recognition, an image is considered valid if it can be recognized as a digit by a human expert, despite potential imperfections. We, therefore, aim to compare in this research question the proportions of valid inputs that DiffGAN and DRfuzz generate. This comparison is crucial because generating a large number of triggering inputs is only beneficial if these inputs are to a large extent valid, as defined above.
@@
We aim in this research question to determine whether the triggering inputs generated by DiffGAN can more effectively guide the training of a model selection mechanism, allowing it to learn the conditions under which one DNN model outperforms another. In many real-world scenarios, it is common to encounter multiple DNNs that perform similarly in terms of overall accuracy, but exhibit significant variations in their performance under specific input conditions [51, 52, 10, 11]. This presents a challenge for developing a model selection mechanism that can dynamically choose the most suitable model based on the input characteristics without having to run all models in real-time. An effective solution to this problem is to use an ML-based selection mechanism, which can predict which of the available models will perform better for a given input [51, 52]. However, training such an ML-based model selection mechanism requires diverse and informative training data that captures and reveals a wide range of behavioural differences between the models. This is where the generation of triggering inputs becomes critical for the selection mechanism to more effectively learn under which condition one model is preferable over the other. Indeed, the effectiveness of the selection mechanism heavily depends on the quality of the triggering inputs used for training. Therefore, we aim to explore in this RQ whether DiffGAN can better guide the training process compared to DRfuzz. It is important to clarify that we do not intend to propose a novel ensemble method or voting mechanism in this work. Instead, this research question focuses on investigating an important application scenario, that is, studying whether DiffGAN helps learn which model fares better under various input conditions and thus trains a model selection mechanism.
@@ 4.3 Datasets and Models
We consider different combinations of datasets and models in our empirical evaluation. More specifically, we consider MNIST [54] and Cifar-10 [55], two well-known image recognition datasets. As shown in Figure 2, the MNIST dataset comprises 70,000 black-and-white images of handwritten digits, divided into a training set of 60,000 images and a test set of 10,000 images. The Cifar-10 dataset comprises 60,000 color images distributed across ten categories (Figure 3), including 50,000 images designated for training and 10,000 images for testing. These datasets were employed in conjunction with 16 different DNN models, which included both custom-developed and widely used architectures such as ResNet, LeNet-5, AlexNet, and inception models [56, 57, 58]. Comprehensive details about the combinations of datasets and models we used, along with their accuracy, are outlined Tables I and II.
As shown in the tables, our experimental study spanned four distinct and common sources of variation per dataset across DNN models, including the model structures, training data, optimizers, and hyperparameters. Our experiments thus cover 16 different DNN models, each selected to evaluate the effectiveness of our test generation approach. For each model pair, the differences in DNN models are highlighted in bold in Tables I and II. Typical differences across models include:
@@ 4.4 Choice of the Fitness Functions
Due to the computational expense of our experiments, we conducted this study on two model pairs with the Cifar-10 dataset. Seeing significant differences across configurations on such pairs would, however, be enough to justify the combined use of divergence and diversity. We studied both the total number of triggering inputs generated by each configuration and their diversity. To quantify diversity, we leveraged two metrics: Shannon Diversity [59, 60] and Geometric Diversity [1]. A detailed explanation of these metrics, along with our methodology for computing diversity scores, is provided in Section 5.3. The results of this ablation study are summarized in Table III.
@@ 5 Empirical Results
In this section, we describe our experimental evaluation and report the results. We first investigate whether we generate more triggering inputs than the baseline with the same testing budget. We then study the validity and diversity of the generated triggering inputs by both approaches. Finally, we present an important application scenario of the generated triggering inputs, namely the development of a machine learning-based mechanism to predict the best model’s output based on input characteristics, and report the corresponding results.
@@
To answer this research question, we executed DiffGAN and DRfuzz considering the same execution time $T\in[1h,2h,3h]$, across four pairs of DNN models and two distinct datasets, resulting in eight unique combinations (Section 4.3). Because of the randomness in DiffGAN and DRfuzz, we re-executed each of them three times and reported the corresponding number of triggering inputs in Tables IV and V. This decision to limit the execution of the approaches to three iterations was initially due to the considerable execution time associated with the experiments. More specifically, our experiments take 96h (8 model pairs x 6h x 2 approaches) if we execute each approach once. As shown in the tables, DiffGAN consistently outperforms DRfuzz in terms of generating more triggering inputs across all models and datasets. More specifically, our results show that DiffGAN generates on average four times more triggering inputs than DRfuzz when considering the same testing budget. We also calculated the Coefficient of Variation (CV) [61], a measure of relative variability that compares the standard deviation to the mean, independent of the data’s range. We computed the CV score for each model pair and testing budget across the three runs. A lower CV indicates higher consistency across different runs. For example, in our analysis of the MNIST dataset, DRfuzz showed a CV between 0.33% and 21.23%,while DiffGAN exhibited a CV between 1.33% and 12.20%, suggesting less variability in the number of generated triggering inputs across the different runs. These observations are also consistent for the Cifar-10 dataset. These CV values demonstrate the relative stability and reliability of DiffGAN in generating triggering images, with less variability compared to DRfuzz.
@@
Table V: RQ1 results: Number of generated triggering images for models trained on the Cifar-10 dataset using DiffGAN and DRfuzz. The #TI (Initial) column indicates the number of triggering inputs identified in the original test dataset for each set of models.
@@
To systematically compare DiffGAN and DRfuzz in terms of generating valid triggering inputs within a given time budget, we considered the inputs generated in RQ1 by both methods. More specifically, we considered all four model pairs using the MNIST dataset and two model pairs for the CIFAR-10 dataset. We ran both approaches for two hours on these models. We did not consider all model pairs in this RQ, as the validity assessments require significant manual effort, particularly for labeling generated inputs to ensure they are contextually valid. The lower resolution and greater complexity of CIFAR-10 images further complicate this process, as accurately labeling and validating inputs in CIFAR-10 is considerably more challenging compared to MNIST. Given these challenges, we opted to limit our validity study to the MNIST dataset to ensure its feasibility.
@@
Manual validity checking. From the non-redundant set, we randomly sampled images for validation, according to the sample size determined in the previous step. To ensure unbiased evaluation, we shuffled images generated by both DiffGAN and DRfuzz and presented them anonymously to three independent expert evaluators. These experts were tasked with assessing the validity of each image based on predefined criteria. Specifically, they labeled images as either valid or invalid. For example for MNIST dataset, invalid images were defined as those that displayed two numbers, no number, or other issues such as (1) Images that were a 50/50 blend of two digits and could not be clearly identified as a specific digit, (2) Incomplete or blended digits that could not be classified into one category by consensus, (3) Images containing random shapes or non-digit drawings, or (4) Images that were completely white or dark, with no discernible content. Examples of valid and invalid inputs generated by both approaches are depicted in Figure 4. To determine the validity of each image, we considered the agreement among the evaluators. If at least two out of three evaluators labeled an image as valid, we considered it valid. To quantify the level of agreement between the evaluators, we employ Cohen’s kappa [63]—a statistical measure that accounts for agreement occurring by chance. In our study, the value of Cohen’s kappa is 70%, indicating a high level of agreement among the evaluators and underscoring the reliability of the manual labeling process.
@@
The metric computes the sum of the proportions of each category multiplied by the logarithm of those proportions. The negative sign ensures that the index is positive, and higher diversity translates into higher values of $H$. To calculate the Shannon diversity for a set of images, each image is first flattened into a one-dimensional array of pixel values. These flattened arrays can then be combined into a single array that represents the distribution of pixel values across the entire image set. The Shannon score, also known as entropy, is then calculated using this distribution, which quantifies the diversity of pixel intensities within the dataset. A higher Shannon score indicates greater diversity within the set of images.
In some contexts, computing the exponential of the Shannon diversity score is used to provide a more intuitive interpretation of the index. Specifically, by calculating $e^{H}$, where $H$ is the Shannon diversity of a set of images, we can derive the effective number of image categories within the set. This value represents the number of equally likely categories that would produce the same level of entropy as observed in the dataset. For example, if $H$ increases from 11 to 12, the exponential value $e^{12}$ is substantially larger (approximately 162,754.79) when compared to $e^{11}$ = 59,874.14, indicating that the dataset is considerably more diverse. This approach allows for the translation of entropy into a more intuitive measure of diversity, making it easier to understand and compare the diversity of different datasets.
@@
In this research question, we employed both the Shannon and Geometric diversity metrics to analyze the dataset, aiming to evaluate image diversity from two complementary perspectives. The Shannon diversity metric offers a probabilistic view, emphasizing the distribution of pixel values. In contrast, the Geometric diversity metric assesses diversity based on spatial relationships and geometric patterns within the image features extracted by VGG16. By integrating these two metrics, we achieve a more comprehensive understanding of the datasets’ diversity.
@@
Models Method
Only Model A is correct
Only Model B is correct
Both models are incorrect
@@
When trained adequately with triggering inputs, the RF model should learn when each classifier is more accurate, effectively serving as a decision mechanism for selecting the optimal classifier for a given input. We choose RF because: (1) it scales effectively to handle a large number of features, and (2) its robustness against overfitting [71, 72, 73]. We do not intend to provide a novel voting mechanism here, but instead, provide an example of an important application scenario for the triggering inputs we generate. It also serves as a basis to compare the value of the inputs generated by DiffGAN to enable such model selection and compare it with that of DRfuzz.